OpenStudy (anonymous):

Determine n between 0 and 8 such that 1111^2222 = n mod 20

OpenStudy (anonymous):

take the help of log

OpenStudy (anonymous):

i would start by reducing 1111 to 11

OpenStudy (anonymous):

help of log? I tried to log both sides and could do that with the left side but ran into problems with the right.

OpenStudy (amistre64):

since the gcd of 1111 and 20 is 1, they are relatviely prime; so we can determine the value of phi(20) to reduce the exponent with

OpenStudy (amistre64):

since \[a^{\phi(n)}\cong1(mod~n)\]when (a,n)=1

OpenStudy (amistre64):

phi(20) counts all the residue classes that are relatively prime to 20, mod20 ... phi(20) = 8 8 divides 2222 how many times? with what left over?

OpenStudy (anonymous):

277 times with 0.75 left over. So multiply 0.75 times 9?

OpenStudy (amistre64):

you keep confusing remainder with decimal

OpenStudy (amistre64):

2222 mod 8 = 6 since 2222 = 8(277) + 6

OpenStudy (anonymous):

Oh I see...by doing long division?

OpenStudy (amistre64):

yes, that is one way to accomplish it

OpenStudy (anonymous):

Okay, I understand that. Thank you for clarifying!

OpenStudy (amistre64):

the other way is to divide 2222 by 8 and ignore the decimal; then you know that 8(277) is close to 2222; and see that you have to add 6 to get all the way there

OpenStudy (amistre64):

so, \[1111^{2222}=n~(20)\] \[1111^{\phi(20)(277)+6}=n~(20)\] \[1111^{\phi(20)(277)}*1111^6=n~(20)\] \[1*1111^6=n~(20)\] \[1111^6=n~(20)\] that at least makes it less daunting

OpenStudy (amistre64):

so, as satellite suggested, reduce 1111; what is 1111 mod 20?

OpenStudy (anonymous):

It would be 55 R 11 Right?

OpenStudy (amistre64):

looks good, and its the R that we want: so 1111 mod 20 = 11 notice that we have 1111^6 1111.1111.1111.1111.1111.1111 = n (20) we can replace all the 1111 with their 11 equals 11.11.11.11.11.11 = n (20) 11^6 = n (20) getting a little easier to read yet?

OpenStudy (anonymous):

Much! So basically now we can solve the problem in the regular way since it is in small enough of a form?

OpenStudy (amistre64):

yes, now it reduced to a workable form

OpenStudy (anonymous):

Would the answer be 88578 R1. So 1?

OpenStudy (amistre64):

id have to work it out to be sure :)

OpenStudy (amistre64):

(11.11).(11.11).(11.11) = n (20) 121.121.121 = n (20) 121 = 20(6) + 1; R 1 1.1.1 = n (20) 1 = n (20) i believe it is, but id better chk the wolf to be certain

OpenStudy (anonymous):

Okay that sounds good! You are a huge help! This problem was rather daunting to me :) You put it in more understandable terms for sure!

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=1111%5E%282222%29%3Dn%28mod20%29 we either did something spectacularly right, or just got lucky ;) n=1

OpenStudy (anonymous):

Wow! Well either way, I'm perfectly content :) Thanks again for your help! It was a real blessing!

OpenStudy (amistre64):

youre welcome, i took number theory this term and its almost over; so that was fun ;) good luck

OpenStudy (anonymous):

You too!! :)